OpenStudy (anonymous):

Let F(x) be an antiderivative of sin^3 (x). If F(1)=0, find F(8).

ganeshie8 (ganeshie8):

\(\large F(8) - F(1) = \int \limits_1^8 \sin^3x dx\)

OpenStudy (anonymous):

So I just find that on my calculator, right?

ganeshie8 (ganeshie8):

evaluate the right side integral

ganeshie8 (ganeshie8):

use substitution ?

OpenStudy (anonymous):

But this is multiple choice question and all the answers are in decimal form, so should I find the answer on my calculator?

ganeshie8 (ganeshie8):

u may use wolfram..

ganeshie8 (ganeshie8):

\(\large F(8) - F(1) = \int \limits_1^8 \sin^3x dx \) \(\large F(8) - F(1) = 0.6322 \) \(\large F(8) - 0 = 0.6322 \) \(\large F(8) = 0.6322 \)

OpenStudy (anonymous):

Thanks a lot.

ganeshie8 (ganeshie8):

np :)